### v1

@@ Abstract
Compositional actions consist of dynamic (verbs) and static (objects) concepts. Humans can easily recognize unseen compositions using the learned concepts. For machines, solving such a problem requires a model to recognize unseen actions composed of previously observed verbs and objects, thus requiring, so-called, compositional generalization ability. To facilitate this research, we propose a novel Zero-Shot Compositional Action Recognition (ZS-CAR) task. For evaluating the task, we construct a new benchmark, Something-composition (Sth-com), based on the widely used Something-Something V2 dataset. We also propose a novel Component-to-Composition (C2C) learning method to solve the new ZS-CAR task. C2C includes an independent component learning module and a composition inference module. Last, we devise an enhanced training strategy to address the challenges of component variation between seen and unseen compositions and to handle the subtle balance between learning seen and unseen actions. The experimental results demonstrate that the proposed framework significantly surpasses the existing compositional generalization methods and sets a new state-of-the-art. The new Sth-com benchmark and code are available at https://github.com/RongchangLi/ZSCAR_C2C.
@@ 2 Related work
Video action recognition aims to identify actions in videos. The main challenge is extracting discriminative action features from videos. Previous CNN-based methods address this challenge in two primary ways: using 2D CNNs with separate temporal operations [58, 35, 31, 30, 32, 34, 27, 28, 57], or jointly using 3D CNNs [23, 56, 60, 19, 5, 13, 12] to dig out spatiotemporal features. Recently, Transformer-based models transform a video into 3D tokens and model the spatiotemporal correlations through attention layers [2, 46, 1, 11, 37, 29, 49], or adapt the visual image encoder in large vision-language models to the video modality [47, 64, 48].
@@ 4.2 From component to composition
In this module, we infer the action (composition) score $\mathbf{s}_{a}\in\mathbb{R}^{{N_{a}}}$ using the prototypes and the verbs and objects (components) scores. We define the conditional scores $\mathbf{S}_{o|v}\in\mathbb{R}^{{N_{v}}\times{N_{o}}}$ and $\mathbf{S}_{v|o}\in\mathbb{R}^{{N_{o}}\times{N_{v}}}$, that represents the confidence of one component given the visual input and another component as the conditions. Accordingly, for one action $a_{i}=(v_{l},o_{k})$, its confidence score can be calculated by either $s_{a,i}=s_{v,l}\cdot s_{o=o_{k}|v=v_{l}}$ or $s_{a,i}=s_{o,k}\cdot s_{v=v_{l}|o=o_{k}}$. The first one relies on the verbs to infer actions, called the dynamics path. Similarly, the other one is called the static path. As shown in Fig. 2, in the dynamics path, we first encode $\mathbf{F}_{x}$ using a static encoder to provide the visual reference for the object component. Then we concatenate the output with the verb prototype $\mathbf{e}_{v,l}$ and fuse them to the compound prototype-visual feature $\mathbf{f}_{v-x,l}$, using an MLP layer. Then we linearly map the cosine similarity $cos(\mathbf{f}_{v-x,l},\mathbf{e}_{o,k})$ to the range of $[0,1]$, obtaining the conditional object score $s_{o=o_{k}|v=v_{l}}$. The static path follows a similar process except for encoding $\mathbf{F}_{x}$ by the dynamics encoder. Last, we average the inference results from the two paths to obtain the final inference output. Take $s_{a,i}$ as the score of the groudtruth action $a_{i}$, the entire model is optimized using:
@@ 4.3 Enhanced training strategy
Risk in generalizing components. Naturally, there is a high probability of exhibiting different component appearances between training and test domains, namely component domain variation. As shown in Fig. 3, the component ‘book’ exhibits notable distinctions between ‘close a book’ and ‘bend a book’. Besides, given the combination of ‘close’ and ‘book’ in the training set, there is a tendency that the model might align features associated with ‘close’ to ‘book’, or vice versa. These dependencies may confuse the model when encountering unseen action categories.
@@
A concern is that the new compositions $\bar{a_{1}}$ or $\bar{a_{2}}$ may be counterfactual. However, it only introduces noise to the component-to-composition module as the component labels still match the visual input. In this way, we still reserve these counterfactual labels as they can provide regularization effects that avoid overfitting to seen verb-object relations. Besides, it is also consistent with human imagination which is not only constrained within reasonable compositions.
@@ 5.2 Ablation Study
The enhanced training strategy. In Tab. 6, we demonstrate the effectiveness of our enhanced training strategy. It shows that the performances exhibit consistent improvements as gradually applying the proposed strategy.
@@
Next, we validate the design of balanced learning of actions. With the vanilla version with $L_{ind}$ as the baseline, we report the accuracy of all seen and unseen actions and HM without adding bias in Tab. 8. It shows that adding only $L_{con}$ leads to performance degradation, but the degradation in seen accuracy is minor. Similarly, adding CutMix only results in a slight improvement, but with an obvious degradation in seen accuracy. Upon CutMix, we add either $L_{con}$ or $L_{new}$. The results indicate that $L_{new}$ yields a universal improvement, while $L_{con}$ contributes positively to seen accuracy. Last, when $L_{new}$ and $L_{con}$ are integrated, they both bring improvements on all the metrics, and the gains are more pronounced than adding either one alone. This indicates that our training strategy enables the model to achieve a better balance between seen and unseen actions.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (62106089, 62020106012, 62332008, 62336004), and in part by the Engineering and Physical Sciences Research Council (EPSRC), U.K. (Grant EP/N007743/1, Grant MURI/EPSRC/DSTL, Grant EP/R018456/1, Grant EP
/V002856/1, and Grant EP/T022205/1).

### v2

@@ Abstract
Compositional actions consist of dynamic (verbs) and static (objects) concepts. Humans can easily recognize unseen compositions using the learned concepts. For machines, solving such a problem requires a model to recognize unseen actions composed of previously observed verbs and objects, thus requiring so-called compositional generalization ability. To facilitate this research, we propose a novel Zero-Shot Compositional Action Recognition (ZS-CAR) task. For evaluating the task, we construct a new benchmark, Something-composition (Sth-com), based on the widely used Something-Something V2 dataset. We also propose a novel Component-to-Composition (C2C) learning method to solve the new ZS-CAR task. C2C includes an independent component learning module and a composition inference module. Last, we devise an enhanced training strategy to address the challenges of component variations between seen and unseen compositions and to handle the subtle balance between learning seen and unseen actions. The experimental results demonstrate that the proposed framework significantly surpasses the existing compositional generalization methods and sets a new state-of-the-art. The new Sth-com benchmark and code are available at https://github.com/RongchangLi/ZSCAR_C2C.
@@ 2 Related work
Video action recognition aims to identify actions in videos. The main challenge is extracting discriminative action features from videos. Previous CNN-based methods address this challenge in two primary ways: using 2D CNNs with separate temporal operations [59, 35, 31, 30, 32, 34, 27, 62, 28, 58], or jointly using 3D CNNs [23, 57, 61, 19, 51, 5, 13, 12] to dig out spatiotemporal features. Recently, Transformer-based models transform a video into 3D tokens and model the spatiotemporal correlations through attention layers [2, 46, 1, 11, 37, 29, 49], or adapt the visual image encoder in large vision-language models to the video modality [47, 66, 48].
@@ 4.2 From component to composition
In this module, we infer the action (composition) score $\mathbf{s}_{a}\in\mathbb{R}^{{N_{a}}}$ using the prototypes and the verbs and objects (components) scores. We define the conditional scores $\mathbf{S}_{o|v}\in\mathbb{R}^{{N_{v}}\times{N_{o}}}$ and $\mathbf{S}_{v|o}\in\mathbb{R}^{{N_{o}}\times{N_{v}}}$, that represents the confidence of one component given the visual input and another component as the conditions. Accordingly, for one action $a_{i}=(v_{l},o_{k})$, its confidence score can be calculated by either $s_{a,i}=s_{v,l}\cdot s_{o=o_{k}|v=v_{l}}$ or $s_{a,i}=s_{o,k}\cdot s_{v=v_{l}|o=o_{k}}$. The first one relies on the verbs to infer actions, called the dynamics path. Similarly, the other one is called the static path. As shown in Fig. 2, in the dynamics path, we first encode $\mathbf{F}_{x}$ using a static encoder to provide the visual reference for the object component. Then we concatenate the output with the verb prototype $\mathbf{e}_{v,l}$ and fuse them to the compound prototype-visual feature $\mathbf{f}_{v-x,l}$, using an MLP layer. Then we linearly map the cosine similarity $cos(\mathbf{f}_{v-x,l},\mathbf{e}_{o,k})$ to the range of $[0,1]$, obtaining the conditional object score $s_{o=o_{k}|v=v_{l}}$. The static path follows a similar process except for encoding $\mathbf{F}_{x}$ by the dynamics encoder. Last, we average the inference results from the two paths to obtain the final inference output. Taking $s_{a,i}$ as the score of the groudtruth action $a_{i}$, the entire model is optimized using:
@@ 4.3 Enhanced training strategy
Risk in generalizing components. Naturally, there is a high probability of exhibiting different component appearances between training and test domains, namely component domain variation. As shown in Fig. 3, the component ‘book’ exhibits a notable difference in appearance between ‘close a book’ and ‘bend a book’. Besides, given the combination of ‘close’ and ‘book’ in the training set, there is a tendency that the model might align features associated with ‘close’ to ‘book’, or vice versa. These dependencies may confuse the model when encountering unseen action categories.
@@
A concern is that the new compositions $\bar{a_{1}}$ or $\bar{a_{2}}$ may be counterfactual. However, it only introduces noise to the component-to-composition module, as the component labels still match the visual input. In this way, we still reserve these counterfactual labels as they can provide regularization effects that avoid overfitting to seen verb-object relations. Besides, it is also consistent with human imagination which is not only constrained within reasonable compositions.
@@ 5.2 Ablation Study
The enhanced training strategy. In Tab. 6, we demonstrate the effectiveness of our enhanced training strategy. It shows that the results exhibit consistent improvements as gradually applying the proposed strategy.
@@
Next, we validate the proposed balanced learning of actions. With the vanilla version with $L_{ind}$ as the baseline, we report the accuracy of all seen and unseen actions and HM without adding bias in Tab. 8. It shows that adding only $L_{con}$ leads to performance degradation, but the degradation in seen accuracy is minor. Similarly, adding just CutMix results in a slight improvement, but with an obvious degradation in seen accuracy. When using CutMix, we add either $L_{con}$ or $L_{new}$. The results indicate that $L_{new}$ yields a universal improvement, while $L_{con}$ contributes positively to seen accuracy. Last, when $L_{new}$ and $L_{con}$ are integrated, they both bring improvements on all the metrics, and the gains are more pronounced than adding either one alone. This indicates that our training strategy enables the model to achieve a better balance between seen and unseen actions.
@@ Acknowledgements
This work is supported in part by the National Key Research and Development Program of China (2023YFF1105102, 2023YFF1105105), the National Natural Science Foundation of China (62020106012, 62332008, 62106089, U1836218, 62336004), the 111 Project of Ministry of Education of China (B12018), the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX22
_2307), and the UK EPSRC (EP/V002856/1, EP/T022205/1).